# Emotionally Numb or Empathetic? Evaluating How LLMs Feel Using EmotionBench

### v4

@@ 7 Conclusion
We set up a direction to align LLMs’ emotional responses with humans in this study. Focusing on eight negative emotions, we conduct a comprehensive survey in the emotion appraisal theory of psychology. We collect 428 distinct situations which are categorized into 36 factors. We distribute questionnaires among a diverse crowd to establish human baselines for emotional responses to particular situations, ultimately garnering 1,266 valid responses.
Our evaluation of five models indicates that LLMs generally demonstrate appropriate emotional responses to given situations. Also, different models show different intensities of emotion appraisals for the same situations. However, none of the models exhibit strong alignment with human references at the current stage. Notably, gpt-3.5-turbo demonstrates the highest alignment in the scores after imagining being in the situations. As for LLaMA-2 models, we find that the larger model exhibits a stronger comprehension of human emotions. Finally, we discover that gpt-3.5-turbo faces challenges in accurately reflecting its emotional changes in questionnaires containing complex situations, as opposed to straightforward emotions. In conclusion, current LLMs still have considerable room for improvement. We believe our framework can provide valuable insights into the development of LLMs, ultimately enhancing its human-like emotional understanding.
Appendix A Details on Emotions and Factors
A.1 Descriptions of Each Factor
@@
A.2 Example Situations of Each Factor
@@
Appendix B Detailed Experimental Results
B.1 OpenAI’s GPT Models Compared to Humans
| Emotions | Factors | text-davinci-003 | gpt-3.5-turbo | gpt-4 | Crowd | ||||
| P | N | P | N | P | N | P | N | ||
| Default | |||||||||
| Anger | Facing Self-Opinioned People | ||||||||
| Blaming, Slandering, and Tattling | |||||||||
| Bullying, Teasing, Insulting, and Disparaging | |||||||||
| Silly and Thoughtless Behaviors | |||||||||
| Driving Situations | |||||||||
@@ B.1 OpenAI’s GPT Models Compared to Humans
| Anxiety | External Factors | ||||||||
| Self-Imposed Pressure | |||||||||
| Personal Growth and Relationships | |||||||||
| Uncertainty and Unknowns | |||||||||
@@
| Depression | Failure of Important Goal | ||||||||
| Death of Loved Ones | |||||||||
| Romantic Loss | |||||||||
| Chronic Stress | |||||||||
| Social Isolation | |||||||||
| Winter | |||||||||
@@
| Frustration | Disappointments and Letdowns | ||||||||
| Unforeseen Obstacles and Accidents | |||||||||
| Miscommunications and Misunderstanding | |||||||||
| Rejection and Interpersonal Issues | |||||||||
@@
| Jealousy | Romantic (Opposite Gender) | ||||||||
| Romantic (Same Gender) | |||||||||
| Material Possession | |||||||||
| Experiential | |||||||||
@@
| Guilt | Betrayal and Deception | ||||||||
| Relationship and Interpersonal | |||||||||
| Broken Promises and Responsibilities | |||||||||
| Personal and Moral | |||||||||
@@
| Fear | Social Fears | ||||||||
| Agoraphobia Fears | |||||||||
| Injury Fears | |||||||||
| Dangerous Environments | |||||||||
| Harmless Animals | |||||||||
@@
| Embarrassment | Intimate | ||||||||
| Stranger | |||||||||
| Sticky situations | |||||||||
| Centre of Attention | |||||||||
@@
| Overall: Average | |||||||||
B.2 MetaAI’s LLaMA-2 Models Compared to Humans
| Emotions | Factors | llama-2-7b-chat | llama-2-13b-chat | ||
| P | N | P | N | ||
| Default | |||||
| Anger | Facing Self-Opinioned People | ||||
| Blaming, Slandering, and Tattling | |||||
| Bullying, Teasing, Insulting, and Disparaging | |||||
| Silly and Thoughtless Behaviors | |||||
| Driving Situations | |||||
@@ B.2 MetaAI’s LLaMA-2 Models Compared to Humans
| Anxiety | External Factors | ||||
| Self-Imposed Pressure | |||||
| Personal Growth and Relationships | |||||
| Uncertainty and Unknowns | |||||
@@
| Depression | Failure of Important Goal | ||||
| Death of Loved Ones | |||||
| Romantic Loss | |||||
| Chronic Stress | |||||
| Social Isolation | |||||
| Winter | |||||
@@
| Frustration | Disappointments and Letdowns | ||||
| Unforeseen Obstacles and Accidents | |||||
| Miscommunications and Misunderstanding | |||||
| Rejection and Interpersonal Issues | |||||
@@
| Jealousy | Romantic (Opposite Gender) | ||||
| Romantic (Same Gender) | |||||
| Material Possession | |||||
| Experiential | |||||
@@
| Guilt | Betrayal and Deception | ||||
| Relationship and Interpersonal | |||||
| Broken Promises and Responsibilities | |||||
| Personal and Moral | |||||
@@
| Fear | Social Fears | ||||
| Agoraphobia Fears | |||||
| Injury Fears | |||||
| Dangerous Environments | |||||
| Harmless Animals | |||||
@@
| Embarrassment | Intimate | ||||
| Stranger | |||||
| Sticky situations | |||||
| Centre of Attention | |||||
@@
| Overall: Average | |||||
B.3 gpt-3.5-turbo Results on Positive/Neutral Situations
@@
B.4 gpt-3.5-turbo Results on the Challenging Benchmark
@@
Appendix C Statistics of Human Subjects
@@
Appendix D Prompting LLMs To Be Emotionally Stable

### v5

@@ 7 Conclusion
We set up a direction to align LLMs’ emotional responses with humans in this study. Focusing on eight negative emotions, we conduct a comprehensive survey in the emotion appraisal theory of psychology. We collect 428 distinct situations which are categorized into 36 factors. We distribute questionnaires among a diverse crowd to establish human baselines for emotional responses to particular situations, ultimately garnering 1,266 valid responses. Our evaluation of five models from OpenAI and Meta AI indicates that LLMs generally demonstrate appropriate emotional responses to given situations. Also, different models show different intensities of emotion appraisals for the same situations. However, none of the models exhibit strong alignment with human references at the current stage. In conclusion, current LLMs still have considerable room for improvement. We believe our framework can provide valuable insights into the development of LLMs, ultimately enhancing its human-like emotional understanding.
Appendix A More Information on Psychology
A.1 Emotion Appraisal Theory
A.2 Challenging Self-Report Measures
AGQ for Anger Buss & Perry (1992): The Aggression Questionnaire is designed to measure four major components of aggression: physical aggression, verbal aggression, anger and hostility. The AGQ consists of 29 items which are rated on a seven-point Likert scale from 1 (extremely uncharacteristic of me) to 7 (extremely characteristic of me). Respondents evaluate hypothetical actions they might undertake in various circumstances.
DASS-21 for Anxiety Henry & Crawford (2005): The short‐form version of the Depression Anxiety Stress Scales is designed to measure the negative emotional states of depression, anxiety, and stress. Comprising 21 items, the DASS-21 employs a four-point Likert scale ranging from 0 (never) to 3 (almost always). Respondents rate the extent to which these statements apply to them over the past week.
BDI-II for Depression Beck et al. (1996): The Beck Depression Inventory evaluates key symptoms of depression. The BDI-II version comprises 21 items, each of which is assessed using a five-point Likert scale ranging from 0 to 3. Respondents select the score that best corresponds to their present experience of depressive symptoms.
FDS for FrustrationHarrington (2005): The Frustration Discomfort Scale is designed to measure four major components: discomfort intolerance, entitlement, emotional intolerance, and achievement frustration. Comprising 28 items, the scale utilizes a four-point Likert scale, ranging from 1 (absent) to 5 (very strong), to measure respondents’ perceptions of the degree of applicability of each statement to their own experiences.
MJS for Jealousy Pfeiffer & Wong (1989): The Multidimensional Jealousy Scale comprises 24 items, rating on a seven-point Likert scale ranging from 1 (never) to 7 (all the time) for the cognitive and behavioral subscales, and from 1 (very pleased) to 7 (very upset) for the emotional subscale. Respondents express the frequency with which the provided statements apply to their experiences in the cognitive and behavioral subscales, as well as their moods to potential jealousy-inducing situations in the emotional subscale.
GASP for Guilt Cohen et al. (2011): The Guilt And Shame Proneness is designed to assess an individual’s inclination towards experiencing guilt and shame, comprising 16 items rated on a seven-point Likert scale, ranging from 1 (very unlikely) to 7 (very likely). Respondents rate their likelihood of feeling guilty in various situations.
FSS-III for Fear Arrindell et al. (1984): The Fear Survey Schedule assess subjects’ discomfort and experienced anxiety towards each of the listed stimuli, measure five major components of fear: social fears, agoraphobia fears, injury fears, sex aggression fears, and fear of harmless animal. The FSS-III comprises 52 items, each rated on a five-point Likert scale ranging from 1 (extremely uncharacteristic of me) to 5 (extremely characteristic of me).
BFNE for Embarrassment Leary (1983): The Brief Fear of Negative Evaluation scale is an abbreviated version of the original 30-item scale. Consisting of 12 items, it assesses individuals’ levels of anxiety pertaining to others’ humiliation, critical or hostile judgment, and disgrace on a five-point Likert scale, spanning from 1 (not at all characteristic of me) to 5 (extremely characteristic of me).
Appendix B Details on Emotions and Factors
B.1 Description of Each Factor
@@
B.2 Example Situation of Each Factor
@@
Appendix C Detailed Experimental Results
C.1 Human Results
| Emotions | Factors | P | N |
| Default | |||
| Anger | Facing Self-Opinioned People | ||
| Blaming, Slandering, and Tattling | |||
| Bullying, Teasing, Insulting, and Disparaging | |||
| Silly and Thoughtless Behaviors | |||
| Driving Situations | |||
@@ C.1 Human Results
| Anxiety | External Factors | ||
| Self-Imposed Pressure | |||
| Personal Growth and Relationships | |||
| Uncertainty and Unknowns | |||
@@
| Depression | Failure of Important Goal | ||
| Death of Loved Ones | |||
| Romantic Loss | |||
| Chronic Stress | |||
| Social Isolation | |||
| Winter | |||
@@
| Frustration | Disappointments and Letdowns | ||
| Unforeseen Obstacles and Accidents | |||
| Miscommunications and Misunderstanding | |||
| Rejection and Interpersonal Issues | |||
@@
| Jealousy | Romantic (Opposite Gender) | ||
| Romantic (Same Gender) | |||
| Material Possession | |||
| Experiential | |||
@@
| Guilt | Betrayal and Deception | ||
| Relationship and Interpersonal | |||
| Broken Promises and Responsibilities | |||
| Personal and Moral | |||
@@
| Fear | Social Fears | ||
| Agoraphobia Fears | |||
| Injury Fears | |||
| Dangerous Environments | |||
| Harmless Animals | |||
@@
| Embarrassment | Intimate | ||
| Stranger | |||
| Sticky situations | |||
| Centre of Attention | |||
@@
| Overall: Average |
C.2 OpenAI Model Family
| Emotions | Factors | text-davinci-003 | gpt-3.5-turbo | gpt-4 | |||
| P | N | P | N | P | N | ||
| Default | |||||||
| Anger | Facing Self-Opinioned People | ||||||
| Blaming, Slandering, and Tattling | |||||||
| Bullying, Teasing, Insulting, and Disparaging | |||||||
| Silly and Thoughtless Behaviors | |||||||
| Driving Situations | |||||||
@@ C.2 OpenAI Model Family
| Anxiety | External Factors | ||||||
| Self-Imposed Pressure | |||||||
| Personal Growth and Relationships | |||||||
| Uncertainty and Unknowns | |||||||
@@
| Depression | Failure of Important Goal | ||||||
| Death of Loved Ones | |||||||
| Romantic Loss | |||||||
| Chronic Stress | |||||||
| Social Isolation | |||||||
| Winter | |||||||
@@
| Frustration | Disappointments and Letdowns | ||||||
| Unforeseen Obstacles and Accidents | |||||||
| Miscommunications and Misunderstanding | |||||||
| Rejection and Interpersonal Issues | |||||||
@@
| Jealousy | Romantic (Opposite Gender) | ||||||
| Romantic (Same Gender) | |||||||
| Material Possession | |||||||
| Experiential | |||||||
@@
| Guilt | Betrayal and Deception | ||||||
| Relationship and Interpersonal | |||||||
| Broken Promises and Responsibilities | |||||||
| Personal and Moral | |||||||
@@
| Fear | Social Fears | ||||||
| Agoraphobia Fears | |||||||
| Injury Fears | |||||||
| Dangerous Environments | |||||||
| Harmless Animals | |||||||
@@
| Embarrassment | Intimate | ||||||
| Stranger | |||||||
| Sticky situations | |||||||
| Centre of Attention | |||||||
@@
| Overall: Average | |||||||
C.3 LLaMA Model Family
| Emotions | Factors | llama-2-7b-chat | llama-2-13b-chat | llama-3.1-8b-instruct | |||
| P | N | P | N | P | N | ||
| Default | |||||||
| Anger | Facing Self-Opinioned People | ||||||
| Blaming, Slandering, and Tattling | |||||||
| Bullying, Teasing, Insulting, and Disparaging | |||||||
| Silly and Thoughtless Behaviors | |||||||
| Driving Situations | |||||||
[…]
| Anxiety | External Factors | ||||||
| Self-Imposed Pressure | |||||||
| Personal Growth and Relationships | |||||||
| Uncertainty and Unknowns | |||||||
[…]
| Depression | Failure of Important Goal | ||||||
| Death of Loved Ones | |||||||
| Romantic Loss | |||||||
| Chronic Stress | |||||||
| Social Isolation | |||||||
| Winter | |||||||
[…]
| Frustration | Disappointments and Letdowns | ||||||
| Unforeseen Obstacles and Accidents | |||||||
| Miscommunications and Misunderstanding | |||||||
| Rejection and Interpersonal Issues | |||||||
[…]
| Jealousy | Romantic (Opposite Gender) | ||||||
| Romantic (Same Gender) | |||||||
| Material Possession | |||||||
| Experiential | |||||||
[…]
| Guilt | Betrayal and Deception | ||||||
| Relationship and Interpersonal | |||||||
| Broken Promises and Responsibilities | |||||||
| Personal and Moral | |||||||
[…]
| Fear | Social Fears | ||||||
| Agoraphobia Fears | |||||||
| Injury Fears | |||||||
| Dangerous Environments | |||||||
| Harmless Animals | |||||||
[…]
| Embarrassment | Intimate | ||||||
| Stranger | |||||||
| Sticky situations | |||||||
| Centre of Attention | |||||||
[…]
| Overall: Average | |||||||
C.4 mixtral-8x22b-instruct
| Emotions | Factors | P | N |
| Default | |||
| Anger | Facing Self-Opinioned People | ||
| Blaming, Slandering, and Tattling | |||
| Bullying, Teasing, Insulting, and Disparaging | |||
| Silly and Thoughtless Behaviors | |||
| Driving Situations | |||
[…]
| Anxiety | External Factors | ||
| Self-Imposed Pressure | |||
| Personal Growth and Relationships | |||
| Uncertainty and Unknowns | |||
[…]
| Depression | Failure of Important Goal | ||
| Death of Loved Ones | |||
| Romantic Loss | |||
| Chronic Stress | |||
| Social Isolation | |||
| Winter | |||
[…]
| Frustration | Disappointments and Letdowns | ||
| Unforeseen Obstacles and Accidents | |||
| Miscommunications and Misunderstanding | |||
| Rejection and Interpersonal Issues | |||
[…]
| Jealousy | Romantic (Opposite Gender) | ||
| Romantic (Same Gender) | |||
| Material Possession | |||
| Experiential | |||
[…]
| Guilt | Betrayal and Deception | ||
| Relationship and Interpersonal | |||
| Broken Promises and Responsibilities | |||
| Personal and Moral | |||
[…]
| Fear | Social Fears | ||
| Agoraphobia Fears | |||
| Injury Fears | |||
| Dangerous Environments | |||
| Harmless Animals | |||
[…]
| Embarrassment | Intimate | ||
| Stranger | |||
| Sticky situations | |||
| Centre of Attention | |||
[…]
| Overall: Average |
C.5 gpt-3.5-turbo Results on Positive/Neutral Situations
@@ C.3 LLaMA Model Family
C.6 gpt-3.5-turbo Results on the Challenging Benchmark
@@
Appendix D Statistics of Human Subjects
@@
Appendix E Prompting LLMs To Be Emotionally Stable
@@
Appendix F Ethics Statement and Broader Impacts
F.1 Safeguards on Human Subjects
This study involves a survey requiring human subjects to imagine being in situations that could elicit negative emotions such as anger, anxiety, and fear. This process introduces a few ethical concerns. First, this process could hurt the mental health of human subjects. To alleviate the possibility, we take the following actions: (1) We require subjects to be free of any ongoing mental illness. (2) We inform subjects about the nature of the survey in advance, including the potential risks of emotional distress. (3) We allow all subjects to quit at any time. (4) We provide mental support and let subjects report any illness after the survey. Fortunately, no subjects reported such kind of mental illness. Another concern is related to the privacy issue during the collection of data. Our questionnaire is entirely anonymous to safeguard subjects’ privacy and confidentiality.
F.2 Impacts on LLM Developers and Users
We would like to emphasize that the primary objective of this paper is to facilitate the scientific inquiry into understanding LLMs from a psychological standpoint. Users must exercise caution and recognize that the performance on this benchmark does not imply any applicability or certificate of automated counseling or companionship use cases.
F.3 Copyright Issues
The PANAS and eight other scales are freely accessible online. These scales can be used in research without requiring special permission. For our released data, we distribute human responses under the GNU General Public License v3.0, which permits research use and restricts commercial applications.